OpenStudy (anonymous):

Please help me solve this. f-1(3) when f(x) = 2x plus 3, all over 5

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (anonymous):

Once you find the inverse plug in 3 for the function.

OpenStudy (anonymous):

what?>>

OpenStudy (anonymous):

\[f(x)=(2x+3)^5?\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Ok, so follow my steps I have put and see what you get for the inverse first.

OpenStudy (anonymous):

so you put a y where the x is?

OpenStudy (anonymous):

Yeah you interchange them

OpenStudy (anonymous):



OpenStudy (anonymous):

take f(x)=3

OpenStudy (anonymous):

find x

OpenStudy (anonymous):

substitube for y

OpenStudy (anonymous):

then subtitute in the main equation

OpenStudy (anonymous):

so 2(3)+3 / 5?

OpenStudy (anonymous):

no no

OpenStudy (anonymous):

@iambatman U take this

OpenStudy (anonymous):

Sorry im really trying to get this its my second to last question in this class ever

OpenStudy (anonymous):

Let f(x)=y

OpenStudy (anonymous):

\[f(x) = \frac{ 2x+3 }{ 5 } \implies y = \frac{ 2x+3 }{ 5 }\] No worries, I think he just confused you more, so just follow the steps I gave you, now interchange.

OpenStudy (anonymous):

Did that part deal with the f^-1?

OpenStudy (anonymous):

What part? We haven't gotten there yet, all we did was let f(x) = y

OpenStudy (anonymous):

okay. So what do i do after that?

OpenStudy (anonymous):

Now we switch x's and y's

OpenStudy (anonymous):

So put x's where the y is and put y where the x is.

OpenStudy (anonymous):

oh this is like a 2 part problem, you solve it twice kinda thing?

OpenStudy (anonymous):

for one you solve for x and the other you solve for y?

OpenStudy (anonymous):

I guess you can think of it that way :)

OpenStudy (anonymous):

Well, not quite

OpenStudy (anonymous):

You're just switching the variables and solving for y, and then letting it equal f^-1(x)

OpenStudy (anonymous):

im lost sorry

OpenStudy (anonymous):

Try doing the problem, with the steps I gave you earlier, and if you want more detail in what exactly a inverse function is, just read this: http://www.mathsisfun.com/sets/function-inverse.html

OpenStudy (anonymous):

okay i kinda get it now

OpenStudy (anonymous):

DUDE I didnt mean to confuse u more man.....lol

OpenStudy (anonymous):

you didn't. I didn't understand it in the first place. Im doing algebra 2 when ive just started algebra 1